Question 15

Given that P(-2 ≤ Z ≤ z) = 0.4019, what is the value of z?

(1) 0.19
(2) 0.19
(3) 0.4247
(4) 0.5753
(5) None of the above
1
Expert's answer
2020-05-05T19:53:56-0400

"P(-2\le Z\le z)=0.4019\\"

then,

"P(-2\le Z\le0)=0.4772"

since "P(-2\le Z\le0)\ge P(-2\le Z\le z)" , z is negative value.

Therefore,

"P(-2\le Z\le0)-P(z\le Z\le0)=0.4019\\
0.4772-P(z\le Z\le0)=0.4019\\
P(z\le Z\le0)=0.0753\\"

that means,

"P(z\ge Z)=0.5-0.0753=0.4247\\
\bold{z=-0.19}"

among those five answers there isn't -0.19 therefore the answer is "5^{th}" one.

(5) none of the above